# Towards Stable and Faithful Inpainting

### 08 Dec 2023 v1

@@ 5.2 Further Analysis of ASUKA
p.t. MAE 22.485 0.799 0.151 11.513 0.354 0.258
f.t. MAE 22.495 0.799 0.150 11.460 0.368 0.256
Table 2: Quantitative comparison of ASUKA using pre-trained MAE (p.t. MAE) from [8] and fine-tuned MAE (f.t. MAE).
Ablation of MAE prior. We compare our fine-tuned MAE with directly adopting the MAE trained in [8] without further fine-tuning. To this end, we train ASUKA with the MAE in [8] using the same training strategy and compare the results in. Results suggest the improvements of fine-tuning MAE, especially on FID and U-IDS. This improvement comes from the better adaptation on the real-world masks.
@@
linear 22.402 0.797 0.155 11.934 0.361 0.227
attn 22.474 0.798 0.152 11.613 0.362 0.234
cross x4 22.488 0.799 0.152 11.762 0.368 0.238
ASUKA 22.495 0.799 0.150 11.460 0.368 0.256
Table 3: Ablation of different alignment modules. Linear adopts linear layer to align feature dimension only; attn adopts a self-attention layer to further align distribution; cross x4 adopts 4 cross-attention layers; while ASUKA adopts 4 self-attention layers.
Ablation of alignment module. Our alignment module includes a dimension and a distribution alignment. We validate the effectiveness of each module by comparing with the following variants step by step: i) linear: Use linear layer to align feature dimension only; ii) attn: Based on i, further use a single self-attention block to align the distribution; iii) cross x4: we instead use learnable query and 4 cross-attention layers to learn the MAE prior. ASUKA adopts 4 self-attention blocks. Results are shown in Table 3. The self-attention block shows improved results on simply aligning the dimension as well as using a cross-attention block to align the distribution. Using 4 self-attention blocks improves the capacity of the alignment module.
@@
KL-VAE [39] 22.454 0.795 0.156 11.949 0.343 0.208
Asy. VAE [64] 22.566 0.801 0.151 11.634 0.361 0.231
color-aug 22.475 0.799 0.152 11.603 0.357 0.229
Ours 22.495 0.799 0.150 11.460 0.368 0.256
Table 4: Quantitative comparison of different decoders for SD.
Ablation of decoder training pipeline. For the decoder, we compare ASUKA with i) KL-VAE: the decoder used in SD; ii) Asy. VAE: the decoder trained by [64]; iii) color-aug: only use color augmentation to train the decoder; Results are shown in Table 4, which shows the superiority of our augmented decoder to ensure the mask-unmask consistency.
@@
Limitation: The “curse” of self-attention. The primary limitation of ASUKA stems from the inefficacy of the MAE prior, mainly due to issues within the self-attention module. Specifically, as showin in Figure 9, the presence of multiple similar objects in an image may lead the MAE to incorrectly predict a similar object in the masked region, conflicting with the goal of object removal. Notably, this curse of self-attention significantly impacts diffusion-based generative models. It results in the inability to accurately follow “blank paper” text prompts, even when employing a substantial classifier-free guidance scale of 9. This issue is not unique to SD but is also a common problem in other advanced text-guided diffusion models, such as OpenAI’s DALL-E 2 [38] and Adobe’s FileFly [1]. Nevertheless, ASUKA has the potential to circumvent this issue by modifying the MAE prior, for instance, by instead using a blank paper image as the input to MAE prior. A more comprehensive solution would involve extra control on self-attention layers in diffusion models, which we leave as future work.
@@ 6 Conclusion
In this paper, we proposed Aligned Stable inpainting with Unknown Areas prior (ASUKA) to achieve stable and faithful inpainting via frozen Stable Diffusion inpainting model (SD). To achieve stability, we adopt a reconstruction-based masked-auto encoder (MAE) as the stable prior for masked region purely from unmasked region. Then we align the stable prior to frozen SD with the proposed alignment module. To achieve fidelity, we resolve the mask-unmask inconsistency of SD in the decoding process. We fine-tune a unmask-region conditioned KL-VAE decoder to perform local harmonization during the decoding process. To validate the efficacy of inpainting algorithms in different image domains and mask types, we introduce an evaluation dataset, named as MISATO, from existing datasets. ASUKA enjoys superior stable and faithful inpainting results than state-of-the-art inpainting algorithms and SD-based variants.
Supplementary Material
7 Extra Analysis
7.1 Comparison on Places2
[…]
Co-Mod [61] 19.413 0.709 0.267 5.794 0.274 0.096
MAT [26] 19.729 0.740 0.202 3.765 0.348 0.195
LaMa [45] 21.503 0.771 0.202 6.693 0.247 0.050
MAE-FAR [8] 22.140 0.779 0.174 3.559 0.307 0.105
SD [39] 20.302 0.746 0.193 1.514 0.375 0.207
SD-text 20.359 0.748 0.191 1.506 0.373 0.202
SD-prompt 20.548 0.753 0.189 1.477 0.390 0.234
ASUKA 20.406 0.752 0.183 1.230 0.413 0.287
Table 5: Quantitative comparison on Places 2.
We evaluation ASUKA and other inpainting algorithms on the standard testing benchmark dataset of Places2 [63], which includes 36500 testing examples. Results on this dataset also validate the efficacy of ASUKA.
7.2 User-Study
Model Consistency (%) Quality (%)
Co-Mod [61] 4.69 5
MAT [26] 7.5 3.33
LaMa [45] 8.44 8.33
MAE-FAR [8] 5.31 5.83
SD [39] 10.63 5.83
SD-text 7.81 15.83
SD-prompt 16.25 15.83
ASUKA 39.38 40
Table 6: User-study of top-1 ratio on consistency and quality among all the inpainting results.
To evaluate the user preference on inpainting algorithms, we conduct an user-study. Specifically, we randomly select 40 testing images along with inpainting results on these images. We ask the user to select the best inpainting results from the following perspectives respectively: i) Consistency: the masked region should be consistent with unmasked region, involving color and texture consistency and not generating new elements; ii) Quality: the fidelity of the inpainting results. The first metric measures the stability while the other measures the fidelity of the inpainting results. We collect 100 valid anonymous questionnaire results, and report the average selection ratio among all the inpainting algorithms. This result validate the efficacy of ASUKA on alignment with human preference.
7.3 The Misalignment Between PSNR, SSIM and Human Preference
The PSNR and SSIM are known to be have bad correlation with human preferences. To show this, we report the PSNR and SSIM scores for the qualitative comparison on MISATO ( Figure 8 in the paper) in Figure 10. Higher PSNR and SSIM do not always indicate better results that human prefers. Particularly, blurred inpainting results usually have higher PSNR and SSIM scores, like the 1st, 2rd, 4th, 8th rows. However, it is clear that blurred results are not expected.
Figure 10: Misalignment between PSNR, SSIM and human preference. Images are same with Figure 8. PSNR and SSIM scores are listed in the bottom, where the best values are marked as red.

### 17 Mar 2024 v2

@@ 5.2 Further Analysis of ASUKA
Ablation of MAE prior. We compare our fine-tuned MAE with directly adopting the MAE trained in [7]. To this end, we train ASUKA with the MAE in [7] using the same training strategy and compare the results in Tab. 2(a). Results suggest the improvements of fine-tuning MAE, especially on FID and U-IDS. This improvement comes from the better adaptation on the real-world masks.
Ablation of alignment module. Our alignment module includes a dimension and a distribution alignment. We validate the effectiveness of each module by comparing with the following variants step by step: i) linear: Use linear layer to align feature dimension only; ii) attn: Based on i, further use a single self-attention block to align the distribution; iii) cross x4: we instead use learnable query and 4 cross-attention layers to learn the MAE prior. ASUKA adopts 4 self-attention blocks. Results are shown in Tab. 2(b). The self-attention block shows improved results compared with only align dimension and cross-attention block. Using 4 self-attention blocks improves the capacity of the alignment module.
Ablation of decoder. For the decoder, we compare ASUKA with i) VAE: the decoder used in SD; ii) + cond.: the decoder conditioned on unmasked image [62]; iii) + color: only use color augmentation to train the decoder; Results are in Tab. 2(d), showing the superiority of our augmented decoder.
Table 2: Ablation stuides of ASUKA.
(a) Comparison of ASUKA using pre-trained (p.t.) MAE v.s. fine-tuned (f.t.) MAE.
@@
p.t. 0.151 11.513 0.354 0.258
f.t. 0.150 11.460 0.368 0.256
(b) Ablation of different alignment modules. Linear adopts linear layer; attn adopts a self-attention layer; cross x4 adopts 4 cross-attention layers; ASUKA adopts 4 self-attention layers.
@@
linear 0.155 11.934 0.361 0.227
attn 0.152 11.613 0.362 0.234
cross x4 0.152 11.762 0.368 0.238
ASUKA 0.150 11.460 0.368 0.256
(c) User-study of top-1 ratio among all the inpainting results. Context-stability (C.S.) measures the coherence between masked region and unmasked surroundings, with a preference of not generating new elements; Visual-consistency (V.C.) measures the color consistency.
Model C.S.(%) V.C.(%)
Co-Mod [59] 3.98 4.98
MAT [25] 7.40 3.20
LaMa [45] 8.18 8.28
MAE-FAR [7] 4.88 5.60
SD [39] 10.58 5.75
SD-text 7.70 15.83
SD-prompt 16.18 15.78
SD-Repaint [32] 1.60 0.55
ASUKA 39.43 40.05
(d) Comparison of different decoders for SD. VAE [39] is the original decoder used by SD; + cond. [62] is the decoder conditioned on unmasked image; + color uses the color augmentation strategy to perform local harmonization task; Ours further combines latent augmentation strategy to handle the gap between generated latent and real latent.
@@
VAE 0.156 11.949 0.343 0.208
+ cond. 0.151 11.634 0.361 0.231
+ color 0.152 11.603 0.357 0.229
Ours 0.150 11.460 0.368 0.256
User-Study. To evaluate the user preference on inpainting algorithms, we conduct an user-study. Specifically, we randomly select 40 testing images. We ask the user to select the best inpainting results from the following perspectives respectively: i) Context-stability (C.S.): the generated region should be context-stable with surrounding unmasked region, with a preference of not generating new elements; ii) Visual-consistency (V.C.) : the visual consistency between masked and unmasked regions. We collect 100 valid anonymous questionnaire results, and report the average selection ratio among all the inpainting algorithms in Tab. 2(c). This result validate the efficacy of ASUKA on alignment with human preference.
@@
Limitation: The "curse" of self-attention. The primary limitation of ASUKA stems from the inefficacy of the MAE prior, mainly due to issues within the self-attention module. Specifically, as shown in Fig. 9, the presence of multiple similar objects in an image may lead the MAE to incorrectly predict a similar object in the masked region, conflicting with the goal of object removal. Notably, this curse of self-attention significantly impacts diffusion-based generative models. It results in the inability to accurately follow "blank paper" text prompts, even when employing a substantial classifier-free guidance scale of 9. This issue is not unique to SD but is also a common problem in other advanced text-guided diffusion models, such as OpenAI’s DALL-E 2 [38] and Adobe’s FileFly [1]. Nevertheless, ASUKA has the potential to circumvent this issue by modifying the MAE prior, for instance, by instead using a blank paper image as the input to MAE prior. A more comprehensive solution would involve extra control on self-attention layers in diffusion models, which we leave as future work.
@@ 6 Conclusion
In this paper, we proposed Aligned Stable inpainting with Unknown Areas prior (ASUKA) to achieve context-stable and visual-consistent inpainting via frozen Stable Diffusion inpainting model (SD). To achieve context-stability, we adopt a reconstruction-based masked auto-encoder (MAE) as the context-stable prior for masked region purely from unmasked region. Then we align the context-stable prior to frozen SD with the proposed alignment module. To achieve visual-consistency, we resolve the mask-unmask color inconsistency of SD in the decoding process. We train an unmask-region conditioned KL-VAE decoder to perform local harmonization during the decoding process. To validate the efficacy of inpainting algorithms in different image domains and mask types, we introduce an evaluation dataset, named as MISATO, from existing datasets. ASUKA enjoys context-stable and visual-consistent inpainting results and superior than state-of-the-art inpainting algorithms and SD-based variants.